22. Assume that the Upto wn Company purchased the bonds of the Antel Company face value of $4,000, a stated annual interest rate of 18%, and These bonds have a total interest payment dates of September 1 and March 1. Uptown purchased these bonds on July 1 at 96 plus four months of accrued interest and a brokerage fee of $50. The entry to record the receipt of interest on September 1 on Uptowns books would be A. Cash Bond Interest Income Bond Interest Receivable Bond Interest Receivable $360 $360 $360 $360 $360 $240 B. Cash C. Cash Bond Interest Income $120 D. Cash $720 Bond Interest Receivable Bond Interest Income $240 $480
Answer #1

In the books bond interest receivable shall have been recorded on purchase.

Bond interest receivable = Face value X rate of interest X Months interest accrued/ 12

= $4,000 X 18% X 4/12 = $240

Interest income for two months from purchase date to interest receipt date is:

Interest income = Face value X rate of interest X Months interest earned/ 12

= $4,000 X 18% X 2/12 = $120

Total interest received = Face value X rate of interest X Months interest paid/ 12

= $4,000 X 18% X 6/12 = $360

Correct option is C